A Church Songbook  Key Index(1)  30 May’96

— Key Index —

Am – The Lord Of Covenant, 3 Am – We Are Marching In Messiah’s Band, 7 C – Far Above All Rule And Authority, 12 C – Garment Of Praise, 25 C – I Was Glad, 26 C – Oh, The Blood, 27 C – The Horse And Rider, 10 C – What A Mighty God We Serve, 1 D – Behold, I Am The Lord, 17 D – Come And Let Us Go, 16 D – Finally, My Brethren, 2 D – Garment Of Praise, 25 D – Hail To The King, 9 D – How Great Is Our God, 23 D – How Lovely Is Your Dwelling Place, 14 D – Nothing But The Blood, 11 D – Oh, The Blood, 27 D – Sing Hallelujah, 13 D – Stand Up And Bless, 18 D – There’s A Light Shining Forth, 22 D – We Will Rejoice, 4 D – What A Mighty God We Serve, 1 E – How Great Is Our God, 23 E – Lift Jesus Higher, 24 E – Oh, The Blood, 27 Em – Praise Him In His Sanctuary, 5 Em – Rise Up, 6 F – Oh, The Blood, 27 G – A Song Of Love, 8 G – I Will Sing Praises, 20 G – Jesus Is The Lord, 19 G – Let God Be Magnified, 21 G – Zephaniah 3:17, 15

Property of a Church  Last Revised: 30 May’96  Material used by permission.